OpenStudy (anonymous):

WDetermine the future value of education: Jenny Franklin estimates that as a result of completing her master’s degree, she will earn $7,000 a year more for the next 40 years. What would be the future value of these additional earnings based on an annual interest rate of 6%?

OpenStudy (anonymous):

Jenny Franklin estimates that as a result of completing her master’s degree, she will earn $7,000 a year more for the next 40 years. a. What would be the total amount of these additional earnings? b. What would be the future value of these additional earnings based on an annual interest rate of 6 percent? (Use Table 1–B in the Chapter Overheads.) Solution: a. $7,000 × 40 = $280,000 b. $7,000 × 154.762 = $1,083,334

OpenStudy (anonymous):

IF you have the same homework I have (Chapter 2) here are the rest of the questions and answers: 2. During a job interview, Pam Thompson is offered a salary of $28,000. The company gives annual raises of 6 percent. What would be Pam’s salary during her fifth year on the job? Solution: Year 1: $28,000 Year 2: $28,000 × 1.06 = $29,680 Year 3: $29,680 × 1.06 = $31,460.80 Year 4: $31,460.80 × 1.06 = $33,348.45 Year 5: $33,348.45 × 1.06 = $35,349.35 (Alternate solution: $28,000 × 1.262 (FV$1 6%, 4 years) = $35,336 3. Calculate the future value of a retirement account in which you deposit $2,000 a year for 30 years with an annual interest rate of 7 percent. (Use the tables in the Chapter overheads.) Solution: $2,000 × 94.461 = $188,922 4. Bill Mason is considering two job offers. Job 1 pays a salary of $36,500 with $4,500 of nontaxable employee benefits. Job 2 pays a salary of $34,700 and $6,120 of nontaxable benefits. Which position would have the higher monetary value? Use a 28 percent tax rate. Solution: Job 1: $36,500 + [$4,500/(1 - 0.28)] = $42,750. Job 2: $34,700 + [$6,120/(1 - 0.28)] = $43,200. 5. Helen Meyer receives a travel allowance of $180 each week from her company for time away from home. If this allowance is taxable and she has a 30 percent income tax rate, what amount will she have to pay in taxes for this employee benefit? Solution: $180 X 52 weeks = $9,360 X 0.30 = $2,808.